Calculate first order partial derivatives of a function "z = x^{2}y - 4x\sqrt{y} - 6y^{2} + 5" at the point M(2;1)


1
Expert's answer
2021-10-15T11:48:55-0400

"z_x=2xy-4\sqrt{y}"

"z_x(2,1)=2\cdot2-4\sqrt{1}=0"


"z_y=x^2-2x/\sqrt{y}-12y"

"z_y(2,1)=2^2-2\cdot2-12=-12"